OpenStudy (18jonea):

A person running a 100 m race is at the starting blocks waiting for the race to begin. At the starters call, the runner starts sprinting. The runner reaches a maximum velocity of 8.9 m/s at the finish line. It takes the runner 12 seconds to reach the finish line. After crossing the finish line, it takes the runner 16 seconds before coming to a complete stop 25 meters later. What is the average acceleration of the runner from the start of the race until the finish line? @irishboy123

OpenStudy (irishboy123):

you go first , mate! your answer is ... ?

OpenStudy (18jonea):

.74166 m/s

OpenStudy (irishboy123):

lol!!

OpenStudy (irishboy123):

let me have a go

OpenStudy (18jonea):

ok

OpenStudy (18jonea):

@IrishBoy123

OpenStudy (irishboy123):

\(\bar a = \dfrac{\Delta v}{\Delta t}\) over the course of the race we can say that: \(\Delta v = 0\) \(\implies \bar a = 0\)

OpenStudy (irishboy123):

where \(\bar a\) is the average acceleration

OpenStudy (18jonea):

8.9/12 @IrishBoy123

OpenStudy (irishboy123):

Nah mate I'd go with gut instinct \(\bar a = \dfrac{\Delta v}{\Delta t}\) \(\Delta v = 0 \implies \bar a = 0\)

OpenStudy (18jonea):

0.74 m/s2 1.6 m/s2 0.56 m/s2 8.3 m/s2 these are my options @IrishBoy123

OpenStudy (18jonea):

?

OpenStudy (18jonea):

@mathmale

OpenStudy (irishboy123):

you can use an Equation of Motion \(x = u t + {1 \over 2}a t^2\) to arrive at \(a = \dfrac {2x}{t^2}\) \( = \dfrac {2(125)}{28^2} = 0.319\)

OpenStudy (irishboy123):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @IrishBoy123 but i still think the proper answer is zero :-)) \(\color{#0cbb34}{\text{End of Quote}}\)